likeabossssssss:

A right rectangular prism has a volume of 2414.68 cm³. The length of the solid is 21.2 cm and the width is 8.5 cm. What is the height of the right rectangular prism? Enter your answer in the box.

likeabossssssss:

@KjSaif

KjSaif:

V = l * w * h 2414.68 = 21.2 * 8.5 * h 2414.68 = 180.2h divide both sides by 180.2 13.4 = h h = 13.4 cm

KjSaif:

Answer : 13.4

likeabossssssss:

thx

KjSaif:

np